# Quantifying Multilingual Performance of Large Language Models Across Languages

### v2

@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4, Claude-3 and LlaMa-3, have demonstrated surprising performance in various NLP tasks Achiam et al. (2023); Ouyang et al. (2022); Touvron et al. (2023); Team et al. (2024); Jiang et al. (2023); Bai et al. (2023). However, the majority of the text datasets are presented in high-resource languages such as English Xie et al. (2024). According to the statistics, for GPT-3 model approximately 92.65% of the training tokens are English and all other languages share the remaining 7.35% training tokens OpenAI (2023). Similarly, English accounts for 89.70% of data for pre-training LlaMa 2 Touvron et al. (2023). This disparity leads to a bias where LLMs exhibit superior performance in high-resource languages like English, German, and French but struggle significantly with low-resource languages.
The imbalance in language representation during the training phase of LLMs means these models are less proficient in low-resource languages. For example, LLM cannot understand the true meaning of some slang terms with specific cultural background, such as Chinese idioms Zhang et al. (2023). Additionally, recent studies have shown that pre-trained models perform poorly in languages with insufficient training data Lankford et al. (2024). These observations highlight the need for a metric that quantitatively assesses LLM performance across different languages, especially for those languages with limited resources.
@@
In this paper, we introduce Language Ranker, a novel method that leverages internal representations to quantitatively evaluate the multilingual capabilities of LLMs, particularly focusing on low-resource languages. We establish the representation of LLMs on the English corpus as a baseline and measure the similarity between this baseline and representations from other languages. This similarity metric serves as the performance score for each language. We validate our approach by applying the Language Ranker to five state-of-the-art open-source LLMs: LlaMa2, LlaMa3, Qwen, Mistral-v0.1, and Gemma. We also compare the Language Ranker’s performance with the proportion of each language in the training datasets and against other established benchmarks. Through comprehensive experiments, our major observations can be summarized as follows:
@@
  • We demonstrate that high-resource languages show higher similarity scores with English, whereas low-resource languages exhibit lower similarity scores, validating the effectiveness of the Language Ranker in measuring language-specific performance.
@@
  • We uncover a strong correlation between the LLM performance and the proportion of languages in the pre-training corpus, providing insights into the impact of training data distribution on LLM performance.
@@
  • We show that high-resource languages are more evenly distributed in the embedding space, while low-resource languages tend to be narrowly clustered. This distribution analysis further supports our intrinsic metric’s reliability in assessing language performance.
2 The Proposed Method
@@
In this section, we will give an introduction to our analysis method. First, we will introduce the dataset that we used in our experiment. Then, we will introduce how to obtain the similarity between English and other languages, as well as how to compare different LLMs’ performances.
2.1 Probing Datasets
@@
We use OPUS-100 Zhang et al. (2020) as our evaluation datasets. OPUS-100 is an English-centric multilingual corpus that covers 100 languages. Each sample consists of text in a non-English language as the original data, with its English translation serving as the target data. For example, {"German": "Ich wollte dir erst noch etwas zeigen.","English": "I wanted to show you something first."}. After filtering, there are 94 subsets containing English, including high-resource languages such as German, French, and Chinese, as well as low-resource languages such as Oriya, Kannada, and Kazakh. Each subset contains 2000 samples.
2.2 Similarity Measurement
@@
We employ cosine similarity to measure the LLMs’ performance gap between the target language and English. Specifically, given two sentences $X=\{x_{i}\}_{i=1}^{n}$ and $Y=\{y_{i}\}_{i=1}^{m}$ representing the text in English and the text in the target language. We use the representation obtained after LLM mapping of the last token $x_{n}$ and $y_{m}$ as the representation of the text and calculate the similarity between them. As we know, LLM consists of several layers of Transformer blocks Vaswani et al. (2017). Therefore, after each layer of mapping by the transformer block, we can get a representation vector $x_{n}^{l}$ and $y_{m}^{l}$, $l=1...H$, where $H$ represents the number of the layer of LLMs. According to Li et al. (2024), the intermediate representation can be briefly summarized by the following equations:
@@
2.3 Rank Correlation Measurement
@@
3 Experiments
In our experiments, we utilize five prominent open-source large models: LlaMa2 Touvron et al. (2023), LlaMa3 Meta-AI (2024), Qwen Bai et al. (2023), Mistral-v0.1 Jiang et al. (2023), and Gemma Team et al. (2024).
We conduct experiments to answer the following research questions: RQ1: Can the Language Ranker effectively quantify the performance of LLMs across multiple languages? (Section 3.1) RQ2: How consistent are the performance rankings of different LLMs when evaluated across a diverse set of languages? (Section 3.2) RQ3: Is the proposed cosine similarity metric correlated with the proportion of a language in the LLMs’ pre-training corpus? (Section 3.3) RQ4: Is the proposed cosine similarity metric correlated with performance on other benchmark tasks for quantifying the multilingual capabilities of LLMs? (Section 3.4)
3.1 Can Language Ranker Quantify LLM Performance Across Languages?
@@
3.2 Comparison Across Different LLMs
@@
(1) Different models display similar results across languages. Figure 1 presents the cosine similarity scores across various layers for four different 7B parameter LLMs: LLaMa, Gemma, Mistral, and Qwen. Four LLMs display a similar trend where high-resource languages (i.e., German, Spanish, French, Malay, and Chinese) consistently exhibit higher cosine similarity scores compared to low-resource languages (i.e., Igbo, Kazakh, Kannada, Oriya, and Turkmen). Figure 2 further corroborates these findings by comparing the rank correlation of the similarity scores across the LLMs. Each LLM’s ranking is used as a baseline, and the remaining three models exhibit ranking patterns that are largely similar to this baseline. This similarity indicates that despite differences in model architecture or training specifics, the relative performance of languages remains consistent across these four models.
(2) Fine-tuning on specific languages will improve its performance. According to the technical report of Qwen Bai et al. (2023), Qwen has additional fine-tuning on the Chinese corpus, which leads to better performance in Chinese. In Figure 1, we observe that for LlaMa2, Gemma, and Mistral, the performance of Chinese is slightly lower than that of other high-resource languages. However, for Qwen, the performance of Chinese is roughly comparable to other high-resource languages and even shows a gradual improvement in the last few layers. This improvement is more clear in Qwen, mainly due to additional fine-tuning of the Qwen model family on the Chinese corpus, as noted in the technical report of Qwen.
Figure 3: Similarity scores of LlaMa2 7B and LlaMa3 8B. From left to right, the languages are Czech (cs), Korean (ko), German (de), Arabic (ar), Italian (it), Bengali (bn), French (fr), Turkish (tr), Turkmen (tk), Hindi (hi), Dutch (nl), Kazakh (kk), Kannada (kn), and Thai (th).
(3) Comparison of LlaMa2 and LlaMa3. We also explore the performance of LlaMa 2 7B and LlaMa 3 8B. From Figure 3, we can observe that in some high-resource languages, such as German, French, and Dutch, both LlaMa 2 and LlaMa 3 perform well, and LlaMa 3 performs better than LlaMa 2, for low-resource languages, both perform poorly, and LlaMa 3 performs worse than LlaMa 2. This phenomenon suggests that there is a certain consistency in the performance of high-resource languages and low-resource languages in similarity metrics, which can be used to distinguish high-resource and low-resource languages.
@@
German 0.17% 0.723 Welsh $\leq 0.01\%$ 0.396
French 0.16% 0.737 Persian $\leq 0.01\%$ 0.300
Swedish 0.15% 0.662 Urdu $\leq 0.01\%$ 0.275
Chinese 0.13% 0.552 Kannada $\leq 0.01\%$ 0.236
@@
LlaMa7B Gemma7B Mistral7B Qwen7B LlaMa7B Gemma7B Mistral7B Qwen7B
@@
Figure 4: Visualization of the embedding space of Gemma-7B for eight languages. Four figures at the top are high-resource languages, Four figures at the bottom are low-resource languages.
3.3 Relationship to Ratio of Training Corpus?
In this section, we explore the relationship between the proportion of each language in the LLaMa2 pre-training corpus and their corresponding performance as measured by the similarity metric. According to the technical report of LlaMa2 Touvron et al. (2023), we obtain the proportion of the pre-training corpus of some languages. Table 1 illustrates this relationship by listing a selection of languages with their proportion in the training data and their similarity scores relative to English. The table is divided into two parts: the left side lists high-resource languages with relatively higher proportions in the LLaMa2 pre-training corpus, and the right side lists low-resource languages with very low proportions ($\leq$ 0.01%). For example, German, with a proportion of 0.17%, has a high similarity score of 0.723, indicating strong performance in comparison to English. This trend suggests that languages with a higher proportion in the pre-training corpus tend to have higher similarity scores, reflecting better model performance. In contrast, low-resource languages like Kannada and Urdu, each with proportions of less than 0.01%, have much lower similarity scores (i.e., 0.236 and 0.275).
@@ 3.3 Relationship to Ratio of Training Corpus?
Table 3: Similarity score of different language pairs of Gemma-7b.
@@
Table 4: Similarity score of different language pairs of llama-7b.
3.4 Correlation with Other Inference Tasks?
@@
We use MLMM-evaluation [^1] https://github.com/nlp-uoregon/mlmm-evaluation as our benchmark dataset to evaluate LLMs’ performances on reasoning tasks in various languages. The benchmark dataset can be used to evaluate the LLM across 26 different languages. It consists of three datasets: ARC Clark et al. (2018), HellaSwag Zellers et al. (2019), and MMLU Hendrycks et al. (2020). We chose ARC and MMLU for evaluation, and both of them are multiple-choice datasets. The ARC dataset consists of 7,787 multiple-choice science questions drawn from a variety of sources. The MMLU dataset contains multiple-choice questions derived from diverse fields of knowledge. We selected five high-resource languages (Chinese, German, French, Spanish, Italian) and five low-resource languages (Kannada, Hindi, Armenian, Marathi, Telugu) for evaluation, randomly selected 100 samples from each language for 4-shot learning prediction, and used accuracy as the metric. The LLMs evaluated are consistent with Figure 1: LlaMa 7B, Gemma 7B, Mistral 7B, and Qwen 7B.
The predicted result is shown in Table 2. From the result, we find that for Gemma, Mistral, and Qwen, the performance of high-resource languages is significantly better than that of low-resource languages, and Gemma performs best. For the LlaMa, the performance in all languages is generally not as good as the first three LLMs. This result shows that LLM reasoning ability in low-resource languages is worse than that in high-resource languages. This result proves that there are differences in performance between high-resource and low-resource languages in reasoning tasks, illustrating the effectiveness of the proposed cosine similarity metric.
4 Further Analysis of Proposed Metric
@@
Building on these insights, in this section, we delve deeper into the proposed metric to explore its credibility and reliability further. Specifically, we aim to answer the following questions: RQ5: Is choosing English as the benchmark a wise choice (Section  4.1)? RQ6: What does the subspace of each language look like (Section  4.2)? RQ7: Is choosing cosine similarity a wise choice (Section  4.3)?
@@
4.1 Why Using English as Baseline?
In the above sections, we choose English as a baseline. This is based on the a priori assumption that low-resource languages generally perform worse than high-resource languages. But if we choose other high-resource languages as baselines, will we get the same performance? In other words, how can we ensure that our metric is not affected by the English language itself? To answer this question, we divided our probing datasets into three types: High Resource-High Resource (H-H), High Resource-Low Resource (H-L), and Low Resource-Low Resource (L-L). To fulfill our requirement, we utilize Tatoeba-Challenge Tiedemann (2020) as our dataset instead of opus-100 because the latter is an English-centric dataset which means there is no Low Resource-Low Resource language pair. Tatoeba-Challenge is a challenge set for machine translation that contains 32G translation units in 2,539 bitexts. The whole data set covers 487 languages linked to each other in 4,024 language pairs. We select four language pairs for each group, English-German (en-de), English-French (en-fr), German-French (de-fr), and Italian-German (it-de) represent H-H; German - Silesian (de-szl), French-Erzya (fr-myv), Italian - Romany (it-ro) and Italian - Uighur (it-uig) represent H-L; Azerbaijani-Turkmen (az-tr), Hungarian-Yiddish (hu-yi), Kabyle-Standard Moroccan Tamazight (kab-SMT) and Mari-Tatar (ma-ta) represent L-L. The results are shown in Table 3 and Table 4.
@@ 4.1 Why Using English as Baseline?
4.2 Deeper Analysis of the Embedding Space
@@
4.3 Why Using Cosine Similarity?
Recent research Steck et al. (2024) has shown that cosine similarity is not always a reliable metric. Inspired by Section 4.2, the quality of the performance of various languages can be clearly judged from the subspace distribution. Therefore, we decided to quantify the performance of these languages from a distribution perspective.
$=\omega^{T}Cov(X)\omega\quad s.t.\quad\omega^{T}\omega=1$
Gemma-7b Mistral-7b
@@ 4.3 Why Using Cosine Similarity?
Table 5: Similarity scores and double variance results for some languages on Gemma-7b and Mistral-7b.
@@
We employ the box plot to show the relationship between the proposed cosine similarity metric and the double variance metric more clearly. From the results of Table 5, we can observe that for each LLM, the languages in the left part are some common high-resource languages, which have higher similarity and lower double variance, while the right part is the opposite for low-resource languages. The second observation is that as the variance increases, the similarity score also tends to decrease. The increase in variance means that the distribution of the subspace becomes uneven and the similarity score decreases accordingly. This shows that the proposed cosine similarity metric can be utilized to roughly measure the quality of distribution of the subspace, which can thus measure the performance of LLM in different languages.
5 Related Work
In this section, we summarize two lines of research that are most relevant to ours.
@@ 5 Related Work
Representation Engineering.  Representation engineering has emerged as an effective approach to enhance the interpretability and transparency of LLMs. Researchers have been leveraging internal representations to tackle various challenges. Zou et al. (2023) summarizes the application of representation engineering in bias, fairness, model editing, and other areas. Schwenk (2007) found that the internal representation of LLM has a certain correlation with time and space, and the internal representation can be employed to represent time and space. Li et al. (2024) found that the representation of the attention head inside the LLM can be used to indicate the correct reasoning direction, probe analysis is further used to correct the internal representation direction to improve the LLM’s performance. Marks and Tegmark (2023) study the structure of LLM representations of true/false statements, proved that language models linearly represent the true/false of factual statements. Ju et al. (2024) used probe technique to detect how LLM stores knowledge layer by layer.
@@
Multilingual Language Model.  Recent research such as Qin et al. (2024) summarizes the recent progress and future trends in multilingual large language models. Ahuja et al. (2024) constructed a benchmark to evaluate LLM’s multilingual ability comprising 22 datasets covering 83 languages. Huang et al. (2023); Qin et al. (2023) have proven that LLM performance varies substantially across different languages and they employ a prompt technique to improve task performance across languages. Wendler et al. (2024) explores how LlaMa works in multilingual tasks and what role English plays in these tasks. The imbalance distribution of training corpus in different languages leads to the bias of LLM towards some high-resource languages such as English Blasi et al. (2021). Some approaches employ multilingual language modeling to alleviate the phenomenon Shen et al. (2024); Kalyan et al. (2021); Conneau et al. (2019). These studies show the importance of strengthening the cross-lingual capabilities of the pre-trained model. Schäfer et al. (2024) found that the presence of a primary language in the training process of LLMs can improve the performance of low-resource languages and lead to a more consistent representation of LLMs in different languages. Liu et al. (2024) found that for English-centric LLMs, although translation into English helps improve the performance of NLP tasks, it is not the best choice for all situations.
6 Conclusions and Future Work
@@
In this work, we propose the Language Ranker to evaluate the performance of LLMs across diverse languages by comparing their internal representations to English. The results show that high-resource languages show higher similarity scores with English, while low-resource languages have lower scores, validating the effectiveness of our metric in assessing language performance. Besides, there is a strong correlation between the performance of LLMs in different languages and the proportion of those languages in the pre-training corpus. Further, results indicate that high-resource languages are more evenly distributed in the embedding space, whereas low-resource languages tend to be narrowly clustered. In the future, we plan to design more comprehensive benchmarks to measure LLM’s capabilities in different languages.
Limitations
The proposed Language Ranker approach provides an initial quantitative way to analyze LLM performance across languages. We acknowledge that the language ranker method we proposed offers only a rough measurement. While our findings indicate a correlation between the similarity scores and the proportion of each language in the pre-training dataset, these scores alone are not sufficient to precisely measure the exact proportions. Our intent was to provide an initial quantitative approach to explore this relationship, and we recognize the need for more comprehensive methods and additional metrics to accurately assess the impact of pre-training data distribution across languages. Furthermore, the method does not account for potential biases or skews that could be present in the multilingual evaluation datasets themselves. The existence of such biases can also introduce noise in the resulting rankings of language abilities for different LLMs.

### v3

# Language Ranker: A Metric for Quantifying LLM Performance Across High and Low-Resource Languages

@@ Introduction
Large Language Models (LLMs), such as GPT-4, Claude-3 and LlaMa-3, have demonstrated remarkable performance in various NLP tasks (Achiam et al. 2023; Ouyang et al. 2022; Touvron et al. 2023; Team et al. 2024; Jiang et al. 2023; Bai et al. 2023). However, this progress masks a critical issue: the stark disparity in LLM performance across languages, particularly disadvantaging low-resource languages common in developing countries. This disparity stems from the overwhelming bias towards high-resource languages, especially English, in training datasets (Xie et al. 2024). For instance, approximately 92.65% of GPT-3’s training tokens are in English, leaving a mere 7.35% for all other languages (OpenAI 2023). Similarly, English accounts for 89.70% of LlaMa2’s pre-training data (Touvron et al. 2023).
This imbalance leads to significant performance gaps between high-resource languages (e.g., English, German, French) and low-resource languages, potentially exacerbating global digital divides. LLMs often struggle with context-specific interpretations in low-resource languages, such as understanding culturally specific expressions or idioms (Zhang et al. 2023). Recent studies confirm that pre-trained models perform poorly in languages with insufficient training data (Lankford, Afli, and Way 2024), highlighting the urgent need for robust methods to quantify and address these disparities. It is urgent to develop a tool to quantify linguistic biases in LLMs, so as to contribute to more equitable AI technologies, potentially improving access and performance for underserved language communities worldwide.
@@
To tackle this challenge, we introduce Language Ranker, a novel metric designed to systematically evaluate LLM performance across diverse languages, with a particular focus on low-resource languages. Our method leverages internal representations of LLMs, establishing English corpus representations as a baseline and measuring the similarity between this baseline and representations from other languages. This similarity score serves as a quantitative measure of language-specific performance. We validate Language Ranker by applying it to five state-of-the-art LLMs: LlaMa2 (Touvron et al. 2023), LlaMa3 (Meta-AI 2024), Qwen (Bai et al. 2023), Mistral-v0.1 (Jiang et al. 2023), and Gemma (Team et al. 2024). Additionally, we analyze the relationship between Language Ranker scores, the proportion of languages in training datasets, and performance on established benchmarks. Our comprehensive experiments yield the following key findings:
@@
  • Experimental results indicate that high-resource languages consistently show higher similarity scores with English, while low-resource languages exhibit lower scores, validating Language Ranker’s effectiveness in quantifying language-specific performance disparities.
@@
  • We uncover a strong correlation between LLM performance and the proportion of languages in pre-training corpora, providing crucial insights into the impact of training data distribution on multilingual capabilities.
@@
  • The analysis of embedding space distributions reveals that high-resource languages are more evenly distributed, while low-resource languages cluster narrowly, further supporting the reliability of the proposed metric.
The Proposed Method
@@
In this section, we will give an introduction to our analysis method. First, we will introduce the dataset that we used. Then, we will introduce how to obtain the similarity between English and other languages, as well as how to compare different LLMs’ performances.  [^1] Our code link: https://github.com/lizh9885/Language-Ranker/
Probing Datasets
@@
We use OPUS-100 (Zhang et al. 2020) as our evaluation datasets. OPUS-100 is an English-centric multilingual corpus that covers 100 languages. Each sample consists of text in a non-English language as the original data, with its English translation serving as the target data. For example, {”German”: ”Ich wollte dir erst noch etwas zeigen.”,”English”: ”I wanted to show you something first.”}. After filtering, there are 94 subsets containing English, including high-resource languages such as German, French, and Chinese, as well as low-resource languages such as Oriya, Kannada, and Kazakh. Each subset contains 2000 samples.
Similarity Measurement
@@
We employ cosine similarity to measure the LLMs’ performance gap between the target language and English. Specifically, given two sentences $X=\{x_{i}\}_{i=1}^{n}$ and $Y=\{y_{i}\}_{i=1}^{m}$ representing the text in English and the text in the target language. Since the current LLMs are all autoregressive models, we use the representation obtained after LLM mapping of the last token $x_{n}$ and $y_{m}$ as the representation of the text and calculate the similarity between them. As we know, LLM consists of several layers of Transformer blocks (Vaswani et al. 2017). Therefore, after each layer of mapping by the transformer block, we can get a representation vector $x_{n}^{l}$ and $y_{m}^{l}$, $l=1...H$, where $H$ represents the number of the layer of LLMs. According to (Li et al. 2024), the intermediate representation can be briefly summarized by the following equations:
@@
Rank Correlation Measurement
@@
Experiments
In our experiments, we utilize five prominent open-source large models: LlaMa2 (Touvron et al. 2023), LlaMa3 (Meta-AI 2024), Qwen (Bai et al. 2023), Mistral-v0.1 (Jiang et al. 2023), and Gemma (Team et al. 2024).
We conduct experiments to answer the following research questions in the following four subsections respectively: RQ1: Can the Language Ranker effectively quantify the performance of LLMs across multiple languages? RQ2: How consistent are the performance rankings of different LLMs when evaluated across a diverse set of languages? RQ3: Is the proposed cosine similarity metric correlated with the proportion of a language in the LLMs’ pre-training corpus? RQ4: Is the proposed cosine similarity metric correlated with performance on other benchmark tasks for quantifying the multilingual capabilities of LLMs?
Can Language Ranker Quantify LLM Performance Across Languages? (RQ1)
@@
Figure 3: Cosine similarity scores across model layers for LlaMa2 13B and Qwen 13B. The graph shows results for the same 10 languages as in Figure 1. The findings reveal that the general trends observed in 7B models persist in these larger models.
Comparison Across Different LLMs (RQ2)
@@
(1) Different models display similar results across languages. Figure 1 presents the cosine similarity scores across various layers for four different 7B parameter LLMs: LlaMa2, Gemma, Mistral, and Qwen. Four LLMs display a similar trend where high-resource languages (i.e., German, Spanish, French, Malay, and Chinese) consistently exhibit higher cosine similarity scores compared to low-resource languages (i.e., Igbo, Kazakh, Kannada, Oriya, and Turkmen). Figure 2 further corroborates these findings by comparing the rank correlation of the similarity scores across the LLMs. Each LLM’s ranking is used as a baseline, and the remaining three models exhibit ranking patterns that are largely similar to this baseline. This similarity indicates that despite differences in model architecture or training specifics, the relative performance of languages remains consistent across these four models.
(2) Fine-tuning on specific languages will improve its performance. According to the technical report of Qwen (Bai et al. 2023), Qwen has additional fine-tuning on the Chinese corpus, which leads to better performance in Chinese. In Figure 1, we observe that for LlaMa2, Gemma, and Mistral, the performance of Chinese is slightly lower than that of other high-resource languages. However, for Qwen, the performance of Chinese is roughly comparable to other high-resource languages and even shows a gradual improvement in the last few layers. This improvement is more clear in Qwen, mainly due to additional fine-tuning of the Qwen model family on the Chinese corpus, as noted in the technical report of Qwen.
(3) Comparison of Larger LLM Sizes. We extended our analysis to explore the performance of the similarity metric in larger language models with 13 billion parameters. Figure 3 presents the results for LlaMa2 13B and Qwen 13B. The findings reveal that the general trends observed in 7B models persist in these larger models. Specifically, the relative performance of high-resource and low-resource languages remains largely unchanged, with a clear separation between the two groups. LlaMa2 13B exhibits more pronounced fluctuations in the initial layers compared to its 7B counterpart, suggesting potentially richer early-stage language representations. These observations highlight both the consistency of language performance patterns across model sizes and the potential for larger models to enhance representations for specific languages.
@@
German 0.17% 0.723 Welsh $\leq$0.010.396
French 0.16% 0.737 Persian $\leq$0.010.300
Swedish 0.15% 0.662 Urdu $\leq$0.010.275
Chinese 0.13% 0.552 Kannada $\leq$0.010.236
@@
LlaMa2 7B Gemma 7B Mistral 7B Qwen 7B LlaMa2 7B Gemma 7B Mistral 7B Qwen 7B
@@
Figure 4: Visualization of the embedding space of Gemma 7B for eight languages. Four figures at the top are high-resource languages, Four figures at the bottom are low-resource languages.
Relationship to Ratio of Training Corpus? (RQ3)
In this section, we explore the relationship between the proportion of each language in the LlaMa2 pre-training corpus and their corresponding performance as measured by the similarity metric. According to the technical report of LlaMa2 (Touvron et al. 2023), we obtain the proportion of the pre-training corpus of some languages. Table 1 illustrates this relationship by listing a selection of languages with their proportion in the training data and their similarity scores relative to English. The table is divided into two parts: the left side lists high-resource languages with relatively higher proportions in the LlaMa2 pre-training corpus, and the right side lists low-resource languages with very low proportions ($\leq$ 0.01%). For example, German, with a proportion of 0.17%, has a high similarity score of 0.723, indicating strong performance in comparison to English. This trend suggests that languages with a higher proportion in the pre-training corpus tend to have higher similarity scores, reflecting better model performance. In contrast, low-resource languages like Kannada and Urdu, each with proportions of less than 0.01%, have much lower similarity scores (i.e., 0.236 and 0.275).
@@
Table 3: Similarity score of different language pairs of Gemma 7B.
@@
Table 4: Similarity score of different language pairs of LlaMa2 7B.
Correlation with Other Inference Tasks? (RQ4)
@@
We use MLMM-evaluation [^2] https://github.com/nlp-uoregon/mlmm-evaluation as our benchmark dataset to evaluate LLMs’ performances on reasoning tasks in various languages. The benchmark dataset can be used to evaluate the LLM across 26 different languages. It consists of three datasets: ARC (Clark et al. 2018), HellaSwag (Zellers et al. 2019), and MMLU (Hendrycks et al. 2020). We chose ARC and MMLU for evaluation, and both of them are multiple-choice datasets. The ARC dataset consists of 7,787 multiple-choice science questions drawn from a variety of sources. The MMLU dataset contains multiple-choice questions derived from diverse fields of knowledge. We selected five high-resource languages (Chinese, German, French, Spanish, Italian) and five low-resource languages (Kannada, Hindi, Armenian, Marathi, Telugu) for evaluation, randomly selected 100 samples from each language for 4-shot learning prediction, and used accuracy as the metric. The LLMs evaluated are consistent with Figure 1: LlaMa2 7B, Gemma 7B, Mistral 7B, and Qwen 7B.
The predicted result is shown in Table 2. From the result, we find that for Gemma, Mistral, and Qwen, the performance of high-resource languages is significantly better than that of low-resource languages, and Gemma performs best. For the LlaMa2, the performance in all languages is generally not as good as the first three LLMs. This result shows that LLM reasoning ability in low-resource languages is worse than that in high-resource languages. This result proves that there are differences in performance between high-resource and low-resource languages in reasoning tasks, illustrating the effectiveness of the proposed cosine similarity metric.
Further Analysis of Proposed Metric
@@
Building on these insights, in this section, we delve deeper into the proposed metric to explore its credibility and reliability further. Specifically, we aim to answer the following questions in the following three subsections: RQ5: Is choosing English as the benchmark a wise choice? RQ6: What does the subspace of each language look like? RQ7: Is choosing cosine similarity a wise choice?
@@
Why Using English as Baseline? (RQ5)
In the above sections, we choose English as a baseline. This is based on the a priori assumption that low-resource languages generally perform worse than high-resource languages. But if we choose other high-resource languages as baselines, will we get the same performance? In other words, how can we ensure that our metric is not affected by the English language itself? To answer this question, we divided our probing datasets into three types: High Resource-High Resource (H-H), High Resource-Low Resource (H-L), and Low Resource-Low Resource (L-L). To fulfill our requirement, we utilize Tatoeba-Challenge (Tiedemann 2020) as our dataset instead of opus-100 because the latter is an English-centric dataset which means there is no Low Resource-Low Resource language pair. Tatoeba-Challenge is a challenge set for machine translation that contains 32G translation units in 2,539 bitexts. The whole data set covers 487 languages linked to each other in 4,024 language pairs. We select four language pairs for each group, English-German (en-de), English-French (en-fr), German-French (de-fr), and Italian-German (it-de) represent H-H; German - Silesian (de-szl), French-Erzya (fr-myv), Italian - Romany (it-ro) and Italian - Uighur (it-uig) represent H-L; Azerbaijani-Turkmen (az-tr), Hungarian-Yiddish (hu-yi), Kabyle-Standard Moroccan Tamazight (kab-SMT) and Mari-Tatar (ma-ta) represent L-L. The results are shown in Table 3 and Table 4.
@@
Deeper Analysis of the Embedding Space (RQ6)
@@
Gemma 7B Mistral 7B
@@
Table 5: Similarity scores and double variance results for some languages on Gemma 7B and Mistral 7B.
Why Using Cosine Similarity? (RQ7)
Recent research (Steck, Ekanadham, and Kallus 2024) has shown that cosine similarity is not always a reliable metric. Inspired by Section Deeper Analysis of the Embedding Space (RQ6), the quality of the performance of various languages can be clearly judged from the subspace distribution. Therefore, we decided to quantify the performance of these languages from a distribution perspective.
$=\omega^{T}Cov(X)\omega\quad s.t.\quad\omega^{T}\omega=1$
@@
We employ the box plot to show the relationship between the proposed cosine similarity metric and the double variance metric more clearly. From Table 5, we can observe that for each LLM, the languages in the left part are some common high-resource languages, which have higher similarity and lower double variance, while the right part is the opposite for low-resource languages. The second observation is that as the variance increases, the similarity score also tends to decrease. The increase in variance means that the distribution of the subspace becomes uneven and the similarity score decreases accordingly. This shows that the proposed cosine similarity metric can be utilized to roughly measure the quality of distribution of the subspace, which can thus measure the performance of LLM in different languages.
Related Work
@@
Representation Engineering.  Representation engineering has emerged as an effective approach to enhance the interpretability and transparency of LLMs. Researchers have been leveraging internal representations to tackle various challenges. Zou et al. (2023) summarizes the application of representation engineering in bias, fairness, model editing, and other areas. Gurnee and Tegmark (2023) found that the internal representation of LLM has a certain correlation with time and space, and the internal representation can be employed to represent time and space. Li et al. (2024) found that the representation of the attention head inside the LLM can be used to indicate the correct reasoning direction, probe analysis is further used to correct the internal representation direction to improve the LLM’s performance. Marks and Tegmark (2023) study the structure of LLM representations of true/false statements, proved that language models linearly represent the true/false of factual statements. Ju et al. (2024) used probe technique to detect how LLM stores knowledge layer by layer.
@@
Multilingual Language Model.  Recent research such as (Qin et al. 2024) summarizes the recent progress and future trends in multilingual large language models. Ahuja et al. (2024) constructed a benchmark to evaluate LLM’s multilingual ability comprising 22 datasets covering 83 languages. Huang et al. (2023); Qin et al. (2023) have proven that LLM performance varies substantially across different languages and they employ a prompt technique to improve task performance across languages. Wendler et al. (2024) explores how LlaMa2 works in multilingual tasks and what role English plays in these tasks. The imbalance distribution of training corpus in different languages leads to the bias of LLM towards some high-resource languages such as English (Blasi, Anastasopoulos, and Neubig 2021). Some approaches employ multilingual language modeling to alleviate the phenomenon (Shen et al. 2024; Kalyan, Rajasekharan, and Sangeetha 2021; Conneau et al. 2019). These studies show the importance of strengthening the cross-lingual capabilities of the pre-trained model. Schäfer et al. (2024) found that the presence of a primary language in the training process of LLMs can improve the performance of low-resource languages and lead to a more consistent representation of LLMs in different languages. Liu et al. (2024) found that for English-centric LLMs, although translation into English helps improve the performance of NLP tasks, it is not the best choice for all situations.
Conclusions and Future Work
@@
In this work, we propose the Language Ranker to evaluate the performance of LLMs across diverse languages by comparing their internal representations to English. The results show that high-resource languages show higher similarity scores with English, while low-resource languages have lower scores, validating the effectiveness of our metric in assessing language performance. Besides, there is a strong correlation between the performance of LLMs in different languages and the proportion of those languages in the pre-training corpus. Further, results indicate that high-resource languages are more evenly distributed in the embedding space, whereas low-resource languages tend to be narrowly clustered. In the future, we plan to design more comprehensive benchmarks to measure LLM’s capabilities in different languages. Besides, we plan to explore the application of the Language Ranker to guide the development of more balanced multilingual training datasets and improve LLM performance on low-resource languages.